# The Role of Privacy Guarantees in Voluntary Donation of Private Data for Altruistic Goals

### v1

@@
Abstract
Voluntary donation of private information for altruistic purposes, such as advancing research, is common. However, concerns about data misuse and leakage may deter individuals from donating their information. While prior research has indicated that Privacy Enhancement Technologies (PETs) can alleviate these concerns, the extent to which these techniques influence willingness to donate data remains unclear.
@@ Abstract
This study conducts a vignette survey ($N=485$) to examine people’s willingness to donate medical data for developing new treatments under four privacy guarantees: data expiration, anonymization, use restriction, and access control. The study explores two mechanisms for verifying these guarantees: self-auditing and expert auditing, and evaluates the impact on two types of data recipient entities: for-profit and non-profit institutions.
Our findings reveal that the type of entity collecting data strongly influences respondents’ privacy expectations, which in part influence their willingness to donate data. Respondents have such high expectations of the privacy provided by non-profit entities that explicitly stating the privacy protections provided makes little adjustment to those expectations. In contrast, statements about privacy bring respondents’ expectations of the privacy provided by for-profit entities nearly in-line with non-profit expectations. We highlight the risks of these respective results as well as the need for future research to better align technical community and end-user perceptions about the effectiveness of auditing PETs and to effectively set expectations about the efficacy of PETs in the face of end-user concerns about data breaches.
1 Introduction
@@
The altruistic analysis of personal health data can bring tremendous societal benefit. For instance, the public health sector could leverage individuals’ medical history and health data to analyze epidemiological patterns, support disease surveillance and risk prediction, diagnose rare and emerging diseases, and improve the development of new treatments [18, 5, 14, 42, 61, 106]. The potential of these applications entails economic benefits: for example, the EU Commission believes that innovations supporting data sharing and data availability may save billions of Euros in the EU health sector [28]. Consequently, the need for personal data in the health sector is ever-increasing [117, 18] — especially with the recent boom of large language models (LLMs), whose capabilities depend on the data they are fed.
Personal health data is already being collected by care providers [14, 61] and personal mobile apps and devices [42]; then, third parties may get access to this data through data sharing agreements that regulate data donation [106]. However, people are often unwilling to donate such sensitive personal data due to privacy concerns [88, 112, 97, 42, 61]. For example, people may be uncomfortable with hospital policies that allow data sharing agreements with third parties without the data owners’ explicit consent [64]. In general, privacy concerns significantly influence (i) people’s adoption of, engagement with, and ability to benefit from technology [15, 89] as well as (ii) the regulations suggested by a country’s legislators [27].
In an effort to foster data collection and analysis while protecting the privacy of data contributors, there has been an increasing focus on deploying privacy-enhancing technologies (PETs) for data storage, data processing, and machine learning (ML) training. For example, encrypted databases such as CryptDB [87] or BlindSeer [81] store user data in encrypted form; hardware-secured platforms [8, 12, 16] additionally rely on Trusted Execution Environments (TEEs) to process data in a secure hardware compartment, so that the confidentiality and integrity of this data (and the code processing it) is guaranteed even during computation; secure Multi-Party Computation (MPC) platforms [44, 25] allow distinct entities to pool their encrypted data and jointly analyze it, without any single entity learning other entities’ data, even indirectly; finally, systems combining all these approaches [119, 32] expand the set of guarantees achievable by any single approach. In addition, many PETs implement differential privacy (DP) [7] to prevent the results of statistical computations on a given sensitive dataset from leaking the presence (or absence) of a specific user’s data.
Despite growing research and deployment of PETs, people feel less control over their data than ever [9]. Arguably, the effort towards strengthening and widening the spectrum of PETs’ privacy-preserving guarantees is insufficient if people’s privacy concerns are not mitigated in practice.
Prior work [30, 2, 6, 10, 31, 33, 3, 118, 120, 75] has sought to meaningfully explain such guarantees—or the technologies that implement them—to the people they aim to protect. We build on this work to evaluate whether such guarantees effectively alleviate the people’s privacy concerns, so that they would be more willing to donate their personal data. While much of the prior work that shares our aim has studied specific, singular technologies or guarantees (often DP [30, 19, 75] or end-to-end encryption (E2EE) [33, 11, 2]), we focus on the more general security and privacy guarantees recurring in PETs:
@@ 1. PG(1):
    Anonymization [74, 96]: data is not linkable to its owner (as defined in [24]);
@@ 2. PG(2):
    Access control [42, 51, 76, 105]: data is accessible only by authorized people, specified in the data collection agreement;
@@ 3. PG(3):
    Data expiration [86, 93, 82]: data is discarded or inaccessible after a given expiration time, specified in the data collection agreement;
@@ 4. PG(4):
    Purpose restriction [32, 20]: data is only used for the stated collection purpose, specified in the data collection agreement;
Additionally, and distinct from prior work, we investigate the role of auditing or verification processes increasingly introduced to produce public auditing results for PETs, specifically:
@@ 1. AG(1):
    Expert auditing [65, 113]: engaging an aggregator-selected external advisory board to audit the system to verify that the PET functions as described in the data collection agreement;
@@ 2. AG(2):
    Self auditing [49, 23]: granting anyone, including donors or external advisors appointed by them, the ability to perform such audits.
Specifically, we investigate the role of these general PETs guarantees and audits of them in shaping privacy expectations and data sharing intentions in the context of medical data donation, asking:
@@ 2. RQ(2):
To address these questions, we conduct a vignette survey ($n=485$) following best practice methodology established by prior work [30, 2, 120]. Each survey respondent is presented with a hypothetical opportunity to donate their health data to help develop a treatment for a specific chronic disease, along with information on how their data will be protected (PG(1)–PG(4) enforced (or not) by AG(1)–AG(2)). Furthermore, we isolate the effect of PG(1)–PG(4) and AG(1)–AG(2) from external confounding factors identified by prior work:
  • Egocentricity [104, 85, 43]: people making decisions depending on personal gain, i.e., whether they (or people close to them) are personally affected by the disease that will benefit from data donation;
@@
  • The data-collection entity [106, 48]: whether the collecting entity is a for-profit or a non-profit organization;
@@
  • Socio-demographics [47, 34, 106]: age, gender, education, donation history, and technical background.
@@
We find that even when told nothing about PETs implemented by the entity, participants are 23% more likely, on average, to expect a non-profit to implement PG(1)–PG(4) and AG(1)–AG(2). As a result of these already high privacy expectations for non-profit organizations, we find that mentioning a specific privacy protection in the survey does not significantly enhance people’s willingness to donate towards non-profit entities: even when no privacy protection is explicitly mentioned, 89% of the participants are willing to donate to a non-profit entity. In contrast, for-profit entities need to effectively prove their privacy protections; indeed, explicitly mentioning privacy protections in the survey does increase privacy expectations of for-profit entities from 50% to the level of non-profit entities. Privacy expectations, in turn, influence willingness to donate.
@@
Furthermore, while the technical community has suggested external audits as a mechanism to enhance trust in PET implementation, our initial inquiry suggests that more work is needed to align expert and end-user perceptions of the effectiveness of such audits. In fact, the effect of audit statements on people’s willingness to donate is limited to a specific scenario involving for-profit entities and auditing to check that purpose restriction (PG(4)) is correctly implemented.
We argue that it is thus essential that non-profit entities correctly implement data privacy, as any future data leak would result in a severe loss of trust. In fact, prior work [4] highlights that users value maintaining expected privacy more than obtaining privacy they did not already expect. In contrast, we highlight the risk of for-profit entities using statements about PETs to engage in ‘privacy washing’ [26] to raise privacy expectations in order to collect data. Simultaneously, we note our respondent’s savviness to the gaps in general PETs guarantees to protect against e.g., data breaches and point to the need for future work to investigate how to effectively set expectations about stronger emerging PETs guarantees in the face of end-user skepticism.
2 Related Work
@@
In this section, we examine prior work on people’s willingness to donate personal data, privacy concerns, the role of privacy-enhancing technologies (PETs), and educational or explanatory strategies related to PETs. Our review primarily focuses on literature specific to personal health data.
@@
Data donation.Data donationData donation. The analysis of personal health data is crucial for medical research, particularly in diagnosing emerging or rare diseases and developing new treatments. Indeed, the COVID-19 pandemic has further intensified the demand for personal health data [94]. Despite the importance of this data, its sensitive nature poses significant obstacles to its collection [92, 21, 22, 71, 17, 35]. Prior work indicates that individuals are generally willing to donate their data for altruistic purposes [98, 5, 84, 5, 95], though there is a notable reluctance when it comes to their health data specifically [70, 72]. However, willingness to donate personal health data increases if the donor or their close family members are directly affected by the disease under investigation [41, 14].
@@
Privacy concerns.Privacy concernsPrivacy concerns. Concerns about the privacy and the misuse of donated data are prominent among potential donors [72]. These concerns include the risk of being identified, discriminated against, or having their personal sensitive data misused or leaked [116, 5, 61, 68]. A common source of these concerns is distrust of the receiving entity [116, 91, 50, 5, 100], often due to fears that the entity might violate consent terms by sharing data with unauthorized third parties [54, 70]. This distrust is intensified when participants are unfamiliar with the recipient entity [116, 91, 50], especially if it is a governmental or for-profit organization [5, 100, 37, 48, 60]. Beyond this intentional data misuse, there are concerns about data leakage caused by hackers or unintentional mishandling [55, 114]. People are also reluctant to interact with entities with any history of data breaches, doubting their ability to safeguard sensitive information [108].
@@
Privacy guarantees (PGs) and auditing.Privacy guarantees (PGs) and auditingPrivacy guarantees (PGs) and auditing. People typically expect the recipient entity to offer some level of privacy which depends on the type of data [55]. Consequently, security experts propose that recipient entities use PETs that implement one or more PGs and an auditing process. We focus on a specific set of 4 PGs: data expiration, data anonymization, use restriction, access control, and auditing (see section 1). In particular, we examine whether offering a given PG influences people’s willingness to donate health data to the recipient entity. Furthermore, we focus on two different auditing processes: expert auditing and self auditing.
To the best of our knowledge, no prior work surveys the effect of auditing (AG(1)–AG(2)) on people’s willingness to donate personal data, and only limited prior work investigates the effect of PG(1)–PG(4). For instance, Leon et al. [62] explored how data access, retention, and usage scope influence willingness to allow behavioral data collection, finding that retention and scope of use impact willingness to share personal information but that people were unwilling to pay more for sites that offered the preferred protections. Valdez and Ziefle [109] used conjoint analysis to study German participants’ willingness to donate health data to different recipients, with different quantified privacy risk, and differing purposes of data collection. They ran two conjoint analysis experiments, one expressing quantified risk of re-identification (for k-anonymity) and the other expressing privacy risk as a combination of sample size and exceptionalism (appropriate for differential privacy). They find similar results across both PETs: that the probability of identification and data recipient entity are the two factors that most influence data donation. Weiss et al. [115] examined participants’ willingness to share sensitive data based on varying epsilon thresholds of differential privacy, finding that the willingness to share is influenced by the purpose of the donation and amount of data subjects held by the entity. Other studies have examined public understanding of data retention and deletion [73], general data anonymization [45], and willingness to donate medical data if it is allowed to be used for secondary purposes [53].
We build upon prior work to present a comprehensive overview of current public attitudes towards mainstream privacy guarantees in health data donation. Specifically, we investigate whether PG(1)–PG(4) and AG(1)–AG(2) influence users’ willingness to donate health data.
@@
Mental models on PETs.Mental models on PETsMental models on PETs. PETs aim to address user privacy concerns. While various PETs are available for health data donation environments [52], prior work indicates that only participants with online privacy literacy tend to have more trusting beliefs in PETs [46]. Studies focusing on sensitive personal data, including health data, reveal similar findings. For instance, Dechand et al. [33] found that most participants do not trust encryption because they do not understand how it works: e.g., many believe that service providers can bypass encryption to gain unauthorized data access. Furthermore, Lerner et al. [63] noted that some participants doubt the existence of true privacy.
Non-experts may misunderstand privacy technologies [118, 78, 33], e.g., viewing encryption merely as restrictive access control [118] or confusing it with data encoding [3]. Additionally, non-expert end users may struggle to understand the consequences of inadequate privacy protections [103, 111, 2, 40, 83, 110, 55]. As a result, even when privacy protections are present, people may have risk expectations that are misaligned with reality. We investigate the alignment between stated privacy guarantees and people’s expectations for how their data will be protected as part of RQ1.
@@
Explaining PETs.Explaining PETsExplaining PETs. To address these misunderstandings, technologists and researchers have worked to explain PETs to the public to address their unfamiliarity with privacy concepts [29]. Prior work includes efforts to explain secure communication, such as end-to-end encrypted messaging, and differential privacy [2, 6, 10, 30, 31, 6, 33, 3, 118, 120, 75]. These efforts remain ongoing, as effectively and scalable setting privacy expectations remains a challenge. Methods found effective to explain PETs include visualizations [101], mental models [99, 66], nutrition labels [59], metaphors [56], short statements [80], and privacy games [107, 57]. In this work we use short statements to briefly explain the relevant guarantees in order to evaluate the impact of these guarantees on participants’ privacy expectations and willingness to donate medical data.
3 Methodology

### v2

# The Role of Privacy Guarantees in Voluntary Donation of Private Health Data for Altruistic Goals

@@
Abstract.
@@
The voluntary donation of private health information for altruistic purposes, such as supporting research advancements, is a common practice. However, concerns about data misuse and leakage may deter people from donating their information. Privacy Enhancement Technologies (PETs) aim to alleviate these concerns and in turn allow for safe and private data sharing. This study conducts a vignette survey ($N=494$) with participants recruited from Prolific to examine the willingness of US-based people to donate medical data for developing new treatments under four general guarantees offered across PETs: data expiration, anonymization, purpose restriction, and access control. The study explores two mechanisms for verifying these guarantees: self-auditing and expert auditing, and controls for the impact of confounds including demographics and two types of data collectors: for-profit and non-profit institutions.
Our findings reveal that respondents hold such high expectations of privacy from non-profit entities a priori that explicitly outlining privacy protections has little impact on their overall perceptions. In contrast, offering privacy guarantees elevates respondents’ expectations of privacy for for-profit entities, bringing them nearly in line with those for non-profit organizations. Further, while the technical community has suggested audits as a mechanism to increase trust in PET guarantees, we observe limited effect from transparency about such audits. We emphasize the risks associated with these findings and underscore the critical need for future interdisciplinary research efforts to bridge the gap between the technical community’s and end-users’ perceptions regarding the effectiveness of auditing PETs.
Privacy Guarantee, Health Data Donation, Privacy Expectation, Auditing
ccs: Security and privacy Social aspects of security and privacy
1. Introduction
@@
The altruistic use of personal health data holds immense potential for societal benefit. For example, the public health sector can leverage individuals’ medical histories and health data to analyze epidemiological trends, enhance disease surveillance and risk prediction, diagnose rare and emerging diseases, and accelerate the development of new treatments (Bietz et al., 2015; Aitken et al., 2016; Bartlett et al., 2018; Greene et al., 2019; Kim et al., 2015; Trinidad et al., 2020). Beyond these applications, there are significant economic advantages. The European Commission estimates that innovations enabling greater data sharing and availability could save billions of euros within the EU health sector (Commission, [n. d.]). As a result, the demand for personal health data continues to grow (Wilbanks and Friend, 2016; Bietz et al., 2015), particularly in light of the recent surge in large language models (LLMs), whose effectiveness heavily depends on the data they are trained on.
Upon the acceptance of data-sharing agreements, personal health data is routinely collected by healthcare providers (Bartlett et al., 2018; Kim et al., 2015) as well as personal mobile apps and devices (Greene et al., 2019). This data is often accessed by third parties through data-sharing agreements that govern data donation practices (Trinidad et al., 2020). However, people frequently self-report that privacy concerns deter them from donating their sensitive information (Redmiles, 2020; Voigt et al., 2020; Silber et al., 2022; Greene et al., 2019; Kim et al., 2015). For instance, many individuals express discomfort with hospital policies that permit data-sharing agreements with third parties without obtaining explicit consent from the data owners (Lewis, 2019).
In an effort to foster data collection and analysis while protecting the privacy of data contributors, there has been an increasing focus on deploying privacy-enhancing technologies (PETs) for data storage (Popa et al., 2011; Pappas et al., 2014), data processing (Arasu et al., 2013; Bajaj and Sion, 2011; Baumann et al., 2015; Gupta et al., 2019; Chowdhury et al., 2019; Al-Rubaie and Chang, 2019; Gaballah et al., 2024), and machine learning (ML) training (Wu et al., [n. d.]; De Viti et al., 2022). For example, CoVault (De Viti et al., 2022) enables multiparty analytics for epidemics-related datasets while preserving data confidentiality, and Anonify (Gaballah et al., 2024) supports decentralized, dual-level anonymity for secure medical data donation.
However, in order for such PETs to impact the availability of data, the guarantees they offer must impact user privacy concerns significantly enough to alter their willingness to donate their data.
In order to evaluate whether the guarantees offered by PETs ultimately impact people’s willingness to donate their data, at least three steps are necessary. First, we must be able to effectively explain guarantees (either general guarantees or those of a particular PET) to the people they aim to protect: end-users. Prior work has focused on doing so for particular PETs, chiefly differential privacy (DP) and end-to-end encryption (E2EE) (Cummings et al., 2021; Abu-Salma et al., 2018; Akgul et al., 2021; Bai, 2019; De Luca et al., 2016; Dechand et al., 2019; Abu-Salma et al., 2017; Wu and Zappala, 2018; Xiong et al., 2020; Nanayakkara et al., 2023). Second, using these explanations, we must evaluate whether the guarantees impact people’s privacy concerns (or expectations), and in turn, their willingness to share data. Prior work has done such evaluations using self-report studies focused on individual PETs, again primarily DP (Cummings et al., 2021; Bullek et al., 2017; Nanayakkara et al., 2023) or E2EE (Dechand et al., 2019; Bai et al., 2020; Abu-Salma et al., 2018)). Third, while rare, we must validate those results in the field to confirm that the effects observed replicate when studying actual vs. intended behavior (e.g., (Dooley et al., 2022b; Acquisti et al., 2013)).
Instead of focusing on a singular PET and it’s accompanying guarantees, we focus on addressing steps one and two for the general security and privacy guarantees that recur across PETs:
@@ 1. PG(1):
    Anonymization (Murthy et al., 2019; Shahid et al., 2022): data is not linkable to its owner (as defined in (Chevrier et al., 2019));
@@ 2. PG(2):
    Access control (Greene et al., 2019; Hu et al., 2006; Ni et al., 2010; Tourani et al., 2017): data is accessible only by authorized people, specified in the data collection agreement;
@@ 3. PG(3):
    Data expiration (Politou et al., 2018; Sarkar et al., 2018; Perlman, 2005): data is discarded or inaccessible after a given expiration time, specified in the data collection agreement;
@@ 4. PG(4):
    Purpose restriction (De Viti et al., 2022; Byun et al., 2005): data is only used for the stated collection purpose, specified in the data collection agreement;
While previous research has explored how the combined presence of guarantees PG(1)-PG(4) affects data-sharing intentions (Aitken et al., 2016; Mori, 2016), our study takes a different approach by focusing on the impact of each guarantee individually. Our objective is to identify which specific protections or factors most significantly influence privacy expectations and, in turn, reported data-sharing intentions.
Additionally, we investigate the role of auditing or verification processes increasingly introduced to produce public auditing results for PETs, specifically:
@@ 1. AG(1):
    Expert auditing (Li et al., 2015; Wang et al., 2011): engaging an aggregator-selected external advisory board to audit the system to verify that the PET functions as described in the data collection agreement;
@@ 2. AG(2):
    Self auditing (Holt and Seamons, 2005; Chen et al., 2022): granting anyone, including donors or external advisors appointed by them, the ability to perform such audits.
We examine the role of these PET guarantees and their audits in shaping privacy expectations and influencing data-sharing intentions within a specific context used in prior work (see e.g., (Cummings et al., 2021; Nanayakkara et al., 2023)): medical data donation. We ask:
@@ 2. RQ(2):
To address these questions, we conduct a vignette survey ($n=494$) following the best practice methodology established by prior work (Cummings et al., 2021; Abu-Salma et al., 2018; Xiong et al., 2020). Each survey respondent is presented with a hypothetical opportunity to donate their health data to help develop a treatment for a specific chronic disease, along with information on how their data will be protected (PG(1)–PG(4) enforced (or not) by AG(1)–AG(2)). Furthermore, we isolate the effect of PG(1)–PG(4) and AG(1)–AG(2) from external confounding factors identified by prior work:
@@
  • The data-collection entity (Trinidad et al., 2020; Hendricks-Sturrup and Lu, 2023): whether the collecting entity is a for-profit or a non-profit organization;
@@
  • Socio-demographics (Hargittai et al., 2020; Dooley et al., 2022a; Trinidad et al., 2020): age, gender, education, donation history, and technical background.
@@
We find that even when told nothing about PETs implemented by the entity, participants are 23% more likely, on average, to expect a non-profit to implement PG(1)–PG(4) and AG(1)–AG(2). As a result of these already high privacy expectations for non-profit organizations, we find that mentioning a specific privacy protection in the survey does not significantly enhance people’s willingness to donate towards non-profit entities: even when no privacy protection is explicitly mentioned, 89% of the participants are willing to donate to a non-profit entity. In contrast, for-profit entities need to effectively demonstrate their privacy protections; indeed, explicitly mentioning privacy protections in the survey does increase privacy expectations of for-profit entities from 50% to the level of non-profit entities. Privacy expectations, in turn, influence the willingness to donate.
@@
Furthermore, while the technical community has suggested external audits as a mechanism to increase trust in PET implementation, our initial inquiry suggests that more work is needed to align the perceptions of the experts and end users about the effectiveness of such audits. In fact, the effect of audit statements on people’s willingness to donate is limited to a specific scenario involving for-profit entities and auditing to check that purpose restriction (PG(4)) is correctly implemented.
We argue that it is critical for non-profit entities to rigorously implement data privacy measures, as any future data leak could lead to a significant loss of trust. Prior research (Acquisti et al., 2013) underscores that users place greater value on maintaining their expected privacy than on gaining additional privacy they did not initially anticipate. In contrast, we highlight the risk of for-profit entities engaging in ”privacy washing” (Cirucci, 2024), where statements about PETs are used to artificially raise privacy expectations to encourage data collection. At the same time, our findings reveal that respondents are perceptive of the limitations in general PET guarantees, particularly concerning protections against data breaches. This underscores the need for future research to explore how to effectively communicate the potential of stronger, emerging PET guarantees and address the skepticism of end-users.
2. Related Work
@@
In this section, we examine prior work on people’s willingness to donate personal data, privacy concerns, the role of privacy-enhancing technologies (PETs), and educational or explanatory strategies related to PETs.
@@
Data donation.Data donationData donation. The analysis of personal health data is crucial for medical research, particularly in diagnosing emerging or rare diseases and developing new treatments. Indeed, the COVID-19 pandemic has further intensified the demand for personal health data (Scott et al., 2020). Despite the importance of this data, its sensitive nature poses significant obstacles to its collection (Fahey and Hino, 2020). Prior work indicates that individuals are generally willing to donate their data for altruistic purposes (Pilgrim and Bohnet-Joschko, 2022; Aitken et al., 2016; Seberger and Patil, 2021), though there is a notable reluctance when it comes to their health data specifically (Middleton et al., 2020; Mori, 2016). However, willingness to donate personal health data increases if the donor or their close family members are directly affected by the disease under investigation (Goodman et al., 2017; Bartlett et al., 2018).
@@
Privacy concerns.Privacy concernsPrivacy concerns. Concerns about the privacy and the misuse of donated data are prominent among potential donors (Mori, 2016). These concerns include the risk of being identified, discriminated against, or having their personal sensitive data misused or leaked (Weitzman et al., 2012; Aitken et al., 2016; Kim et al., 2015; McCormick and Hopkins, 2021). A common source of these concerns is distrust of the receiving entity (Weitzman et al., 2012; Howe et al., 2018; Aitken et al., 2016; Stockdale et al., 2019), often due to fears that the entity might violate consent terms by sharing data with unauthorized third parties (Kacsmar et al., 2022; Middleton et al., 2020). This distrust is intensified when participants are unfamiliar with the recipient entity (Weitzman et al., 2012; Howe et al., 2018), especially if it is a governmental or for-profit organization (Aitken et al., 2016; Stockdale et al., 2019; Garrison et al., 2016; Hendricks-Sturrup and Lu, 2023; Khanra, 2023; American Medical Association, 2022). Beyond this intentional data misuse, there are concerns about data leakage caused by hackers or unintentional mishandling (Seh et al., 2020). People are also reluctant to interact with entities with any history of data breaches, doubting their ability to safeguard sensitive information (Seh et al., 2020).
@@
Mental models on PETs.Mental models on PETsMental models on PETs. PETs aim to address user privacy concerns. While various PETs are available for health data donation environments (Jordan et al., 2022), prior work indicates that participants with greater online privacy literacy tend to have more trusting beliefs in PETs (Harborth and Pape, 2020). Studies focusing on sensitive personal data, including health data, reveal similar findings. For instance, Dechand et al. (Dechand et al., 2019) found that most participants do not trust encryption because they do not understand how it works: e.g., many believe that service providers can bypass encryption to gain unauthorized data access. Furthermore, Lerner et al. (Lerner et al., 2022) noted that some participants doubt the existence of true privacy.
Non-experts may misunderstand privacy technologies (Wu and Zappala, 2018; Oates et al., 2018; Dechand et al., 2019), e.g., viewing encryption merely as restrictive access control (Wu and Zappala, 2018) or confusing it with data encoding (Abu-Salma et al., 2017). Additionally, non-expert end users may struggle to understand the consequences of inadequate privacy protections (Tang et al., 2021; Velykoivanenko et al., 2022; Abu-Salma et al., 2018; Gerber et al., 2019; Pham and Castro, 2019; Vaniea et al., 2014). As a result, even when privacy protections are present, people may have risk expectations that are misaligned with reality. We investigate the alignment between stated privacy guarantees and people’s expectations for how their data will be protected as part of RQ1.
@@
Explaining PETs.Explaining PETsExplaining PETs. To address these misunderstandings and evaluate the impact of PETs on downstream factors such as privacy concerns (or expectations) and willingness to share data, technologists and researchers have worked to explain PETs to the public to address their unfamiliarity with privacy concepts (Cranor et al., 2006). Prior work focuses heavily on E2EE and DP (Abu-Salma et al., 2018; Bai, 2019; Cummings et al., 2021; De Luca et al., 2016; Akgul et al., 2021; Dechand et al., 2019; Abu-Salma et al., 2017; Wu and Zappala, 2018; Xiong et al., 2020; Nanayakkara et al., 2023). These efforts remain ongoing, as effectively and scalable setting privacy expectations remains a challenge. Methods found effective to explain PETs include visualizations (Stransky et al., 2021), mental models (Stewart and Lacey, 2012; Lin et al., 2012), nutrition labels (Kelley et al., 2010), metaphors (Karegar et al., 2022), short statements (Papacharissi and Fernback, 2005), and privacy games (Tupsamudre et al., 2018; Karl and Tao, 2023). In this work we use short statements to briefly explain the relevant guarantees in order to evaluate the impact of these guarantees on participants’ privacy expectations and willingness to donate medical data.
@@
Privacy guarantees (PGs) and auditing.Privacy guarantees (PGs) and auditingPrivacy guarantees (PGs) and auditing. Prior work on the effect of PETs and their accompanying PGs typically focuses on the effect of a combination of guarantees offered by a specific PET (e.g., DP (Valdez and Ziefle, 2019; Cummings et al., 2021) or E2EE (Dechand et al., 2019)) on privacy concerns . A more limited subset of studies evaluate the impact on willingness to share data (Valdez and Ziefle, 2019).
Instead of focusing on a specific PET and it’s combined accompanying PGs we seek to examine and compare four PGs that are shared across many PETs: data expiration, data anonymization, use restriction, access control, and auditing (see § 1). In particular, we examine whether offering a given PG influences people’s willingness to donate health data to the recipient entity. Furthermore, we focus on the effect of two different auditing processes – expert auditing and self auditing – on enhancing privacy expectations and willingness to share data. Other studies have examined public understanding of data retention (Murillo et al., 2018) and general data anonymization (Haddow et al., 2011), and one study examines willingness to donate medical data if it is allowed to be used for secondary purposes (Juga et al., 2021). To our knowledge, no prior work focuses on comparing the effects of the four core PGs we study nor the effects of auditing PETs on user privacy expectations or willingness to donate data.
3. Methodology